# Parton Route to Adiabatic Continuity between Integer and Fractional Quantum Hall Effects

### 13 Jan 2024 v1

@@ I Introduction
While our model is motivated by the parton theory of the FQHE, it is interesting to ask if it can be realized by considering systems where partons are real (as opposed to fictitious) fermions. In principle, it applies to a system of spin-up and spin-down electrons each at an integral filling. In the limit when the electrons are non-interacting, both the up and down spin electrons fill an integer number of LLs, which is the case at even integer fillings in typical experiments. In the other limit with a strong onsite attraction, they exhibit a FQHE state of bosonic bound states consisting of two electrons. Achieving a strong attractive interaction between spin-up and spin-down electrons appears unrealistic, however. The same is true for a system of fermions in two layers. Perhaps a more promising platform is ultra-cold atomic systems. One can, in principle, achieve FQHE states in cold atom systems by creating a synthetic magnetic field through either rotation or complex hopping phases. We refer the reader to Ref. [32] for recent progress as well as for references to the earlier literature. One can thus imagine starting with the IQHE states of two species of noninteracting fermions (which could be the same fermionic atoms but in different internal states). By tuning through the Feshbach resonance it is possible to make the short range interaction between the different species of fermions dominant and attractive. Our study predicts that as the strength of the attractive interaction is enhanced, the system will evolve, without gap closing, into a FQHE state of bosonic molecules. This is a possible way to generate bosonic FQHE states in cold atom systems. The physics here is reminiscent of the BCS to BEC crossover [33, 34], which has been realized in ultra-cold Fermi gases [35, 36, 37].
@@
The spherical geometry  [7], which considers $N$ electrons moving on the surface of a sphere subjected to a radial magnetic field produced by an enclosed magnetic monopole, is very convenient for the study of the FQHE. As noted in the introduction, this geometry is not suitable for the problem at hand, which requires a consideration of arbitrarily high LLs. To circumvent this issue, we begin by constructing a lattice analog of the quantum Hall problem in the spherical geometry [7, 43, 44]. We choose a regular icosahedron because it is the Platonic solid with the most faces [45] resembling a sphere. This similarity is expected to result in a better representation of a continuous sphere than that provided by other polygons. We subdivide each triangle into 4 new equilateral triangles and repeat the process $n$ times [46]. We call the resulting figure an $n$-icosahedron. Figure 1(a) shows illustrations with $n=0,1,2$. As mentioned below, this subdivision structure proves helpful for assigning the Peierls phases that describe the magnetic field. The numbers of faces, edges, and vertices of an $n$-icosahedron, denoted by $F_{n},E_{n}$ and $V_{n}$, respectively, satisfy the recurrence relations $F_{n+1}=4F_{n}$, $E_{n+1}=2E_{n}+3F_{n}$, and $V_{n+1}=V_{n}+E_{n}$. Using $(F_{0},E_{0},V_{0})=(20,30,12)$, we have
@@ IV Numerical diagonalization
Figures 4(b) and 4(d) show the spectrum for systems with one additional flux for each species (or two additional fluxes for the bosonic system). These are the quasihole states. Again, there is no adiabatic continuity, as expected. For Fig. 4(b) there is a 9-fold degeneracy at $U=0$ while a 6-fold degeneracy at $U=20$. The degeneracies in the two limits are consistent with the known degeneracies on sphere. For example, for the $42N=$ 4 particle system: the angular momentum of each quasihole for the $(1,1)$ state is 1, producing $L=1\otimes 1=0\oplus 1\oplus 2$, and for the 1/2 Laughlin state with two quasiholes, we have $L=0\oplus 2$. Figure 4(e) plots the spectrum for quasiparticle states of $62N=$ 6 fermions. The obtained degeneracies at $U=20$ in Fig. 4(e) are consistent with the corresponding degeneracies on sphere.

### 23 Jan 2024 v2

# Crossover from Integer to Fractional Quantum Hall Effect

@@ I Introduction
Figure 1: A schematic depiction of the crossover from the IQHE of two species of fermions to the FQHE of their bosonic molecules. The left panel shows two species of fermions, red and blue, each at filling factor one. Each thick arrow represents a flux of magnitude $h/e$. As the inter-species attractive interaction is dialed up, the fermions form bosonic molecules at filling factor 1/2, because now there are twice as many $h/2e$ flux quanta. The bosonic molecules capture a single $h/2e$ flux quantum each (represented by a thin arrow) to become composite fermions at filling factor one, thereby forming an incompressible FQHE state.
@@
While our model is motivated by the parton theory of the FQHE, it is interesting to ask if it can be realized by considering systems where partons are real (as opposed to fictitious) fermions. In principle, it applies to a system of spin-up and spin-down electrons each at an integral filling. In the limit when the electrons are non-interacting, both the up and down spin electrons fill an integer number of LLs, which is the case at even integer fillings in typical experiments. In the other limit with a strong onsite attraction, they exhibit a FQHE state of bosonic bound states consisting of two electrons. Achieving a strong attractive interaction between spin-up and spin-down electrons appears unrealistic, however. The same is true for a system of fermions in two layers. Perhaps a more promising platform is ultra-cold atomic systems. One can, in principle, achieve FQHE states in cold atom systems by creating a synthetic magnetic field through either rotation or complex hopping phases. We refer the reader to Ref. [32] for recent progress as well as for references to the earlier literature. One can thus imagine starting with the IQHE states of two species of noninteracting fermions (which could be the same fermionic atoms but in different internal states). By tuning through the Feshbach resonance it is possible to make the short range interaction between the different species of fermions dominant and attractive. Our study predicts that as the strength of the attractive interaction is enhanced, the system will evolve, without gap closing, into a FQHE state of bosonic molecules. This is a possible way to generate bosonic FQHE states in cold atom systems. The physics here is reminiscent of the BCS to BEC crossover [33, 34], which has been realized in ultra-cold Fermi gases [35, 36, 37]. Figure 1 illustrates the physics pictorially.
@@
The spherical geometry  [7], which considers $N$ electrons moving on the surface of a sphere subjected to a radial magnetic field produced by an enclosed magnetic monopole, is very convenient for the study of the FQHE. As noted in the introduction, this geometry is not suitable for the problem at hand, which requires a consideration of arbitrarily high LLs. To circumvent this issue, we begin by constructing a lattice analog of the quantum Hall problem in the spherical geometry [7, 43, 44]. We choose a regular icosahedron because it is the Platonic solid with the most faces [45] resembling a sphere. This similarity is expected to result in a better representation of a continuous sphere than that provided by other polygons.
We subdivide each triangle into 4 new equilateral triangles and repeat the process $n$ times [46]. We call the resulting figure an $n$-icosahedron. Figure 2(a) shows illustrations with $n=0,1,2$. As mentioned below, this subdivision structure proves helpful for assigning the Peierls phases that describe the magnetic field. The numbers of faces, edges, and vertices of an $n$-icosahedron, denoted by $F_{n},E_{n}$ and $V_{n}$, respectively, satisfy the recurrence relations $F_{n+1}=4F_{n}$, $E_{n+1}=2E_{n}+3F_{n}$, and $V_{n+1}=V_{n}+E_{n}$. Using $(F_{0},E_{0},V_{0})=(20,30,12)$, we have
@@ IV Numerical diagonalization
Figures 5(b) and 5(d) show the spectrum for systems with one additional flux for each species (or two additional fluxes for the bosonic system). These are the quasihole states. Again, there is no adiabatic continuity, as expected. For Fig. 5(b) there is a 9-fold degeneracy at $U=0$ while a 6-fold degeneracy at $U=20$. The degeneracies in the two limits are consistent with the known degeneracies on sphere. For example, for the $42N=$ 4 particle system: the angular momentum of each quasihole for the $(1,1)$ state is 1, producing $L=1\otimes 1=0\oplus 1\oplus 2$, and for the 1/2 Laughlin state with two quasiholes, we have $L=0\oplus 2$. Figure 5(e) plots the spectrum for quasiparticle states of $62N=$ 6 fermions. The obtained degeneracies at $U=20$ in Fig. 5(e) are consistent with the corresponding degeneracies on sphere.